### 12 Apr 2026 v1

# A remark on comparison of the sum and the maximum of positive random variables

Abstract.
We disprove a conjecture stated in a recent paper by Arnold and Villasenor concerning the sum and the maximum of independent and identically distributed half-normal random variables.
Key words and phrases:
maximum; convolution; half-normal distribution; generalized gamma distribution; subexponential distribution
2020 Mathematics Subject Classification:
60E05; 62E10
1. Introduction
Recently, Arnold and Villasenor [1] provided two proofs that the distributions of the sum and the maximum of two independent half-normal random variables are of the same type. More specifically, it holds that
(1.1) $X_{1}+X_{2}=\sqrt{2}\max\{X_{1},X_{2}\}\quad\textup{ in distribution}$
if $X_{1}$ and $X_{2}$ are independent and follow a common half-normal distribution. They also showed that under some regularity conditions, (1.1) implies that $X_{1}$ follows a half-normal distribution. In [1, Section 4], they conjectured that for $n\geq 3$,
(1.2) $X_{1}+\cdots+X_{n}=(n!)^{1/n}\max\{X_{1},\dots,X_{n}\}\quad\textup{ in distribution}$
if $X_{1},\dots,X_{n}$ are independent and follow a common half-normal distribution.
The purpose of this note is to disprove the conjecture. Our arguments are applicable to a subclass of generalized gamma distributions.
Assume that $X_{i},i\geq 1$, are independent and identically distributed real-valued random variables. Let $S_{n}\coloneqq\sum_{i=1}^{n}X_{i}$ and $M_{n}\coloneqq\max_{1\leq i\leq n}X_{i}$ for $n\geq 1$.
Theorem $1.$
Assume that $X_{1}$ has density
$f(x)=c_{1}\exp\left(-c_{2}x^{\beta}\right),x\geq 0,$
for some constants $c_{1},c_{2}>0$ and $\beta>0$. Assume that $n\geq 2$ and
(1.3) $\beta<\frac{n\log n}{n\log n-\log(n!)}.$
Then there is no constant $C$ such that $S_{n}=CM_{n}$ in distribution.
If $c_{1}=\sqrt{2/\pi},c_{2}=1/2$, and $\beta=2$, then $X_{1}$ follows the standard half-normal distribution, and, (1.3) holds if and only if $n\geq 3$. We see this by induction on $n$, using $n>\exp(\beta-1)$.
We show Theorem 1.1 by contradiction. The main idea of the proof is to first compare the asymptotic behavior of $P(S_{n}\leq x)$ and $P(M_{n}\leq x)$ as $x\to+0$ and then compare the asymptotic behavior of $P(S_{n}>x)$ and $P(M_{n}>x)$ as $x\to+\infty$.
2. Proof
First, we compare behaviors of $P(S_{n}\leq x)$ and $P(M_{n}\leq x)$ as $x\to+0$.
Lemma 2.1.
Assume that $X_{1}$ has a density function $f$ such that $f(0)>0$ and $f$ is continuous at $0$. If there exists a constant $C$ such that $S_{n}=CM_{n}$ in distribution, then $C=(n!)^{1/n}$.
Proof.
We see that for every $x>0$,
$P(S_{n}\leq x)=\int_{D_{1}}\prod_{i=1}^{n}f(x_{i})dx_{1}\cdots dx_{n},$
where $D_{1}\coloneqq\{(x_{1},\dots,x_{n})\in[0,\infty)\colon x_{1}+\cdots+x_{n}\leq x\}$. By the change of variables $x_{i}=xu_{i}$, we obtain that for every $x>0$,
$P(S_{n}\leq x)=x^{n}f(0)^{n}\int_{D_{2}}\prod_{i=1}^{n}\frac{f(xu_{i})}{f(0)}du_{1}\cdots du_{n},$
where $D_{2}\coloneqq\{(u_{1},\dots,u_{n})\in[0,\infty)\colon u_{1}+\cdots+u_{n}\leq 1\}$. Since $f$ is continuous at $0$, we can apply the dominated convergence theorem and obtain that
$\lim_{x\to+0}\int_{D_{2}}\left|\prod_{i=1}^{n}\frac{f(xu_{i})}{f(0)}-1\right|du_{1}\cdots du_{n}=0.$
By this and the equality that $\int_{D_{2}}du_{1}\cdots du_{n}=\frac{1}{n!}$, we obtain that
(2.1) $\lim_{x\to+0}\frac{n!P(S_{n}\leq x)}{x^{n}f(0)^{n}}=1.$
We see that for every $x>0$,
$P(M_{n}\leq x)=P(X_{1}\leq x)^{n}=f(0)^{n}\left(\int_{0}^{x}\frac{f(t)}{f(0)}dt\right)^{n}=f(0)^{n}\left(x+\int_{0}^{x}\frac{f(t)}{f(0)}-1\,dt\right)^{n}.$
Since
$\lim_{x\to+0}\frac{1}{x}\int_{0}^{x}\left|\frac{f(t)}{f(0)}-1\right|\,dt=0,$
we see that
(2.2) $\lim_{x\to+0}\frac{P(M_{n}\leq x)}{x^{n}f(0)^{n}}=1.$
By (2.1) and (2.2), we see that $C=(n!)^{1/n}$. ∎
Second, we compare behaviors of $P(S_{n}>x)$ and $P(M_{n}>x)$ as $x\to+\infty$. By Lemma 2.1, $S_{n}=(n!)^{1/n}M_{n}$ in distribution. Without loss of generality, we can let $c_{2}=1$.
We first consider the case where $\beta\geq 1$.
We give a lower bound for $P(M_{n}>x)$. By the mean value theorem, we see that
$P(M_{n}>x)=1-P(X_{1}\leq x)^{n}\geq nP(X_{1}>x)P(X_{1}\leq x)^{n-1}.$
Since there exists $r_{1}>0$ such that $P(X_{1}\leq r_{1})\geq 1/2$, we see that
$P(M_{n}>x)\geq\frac{n}{2^{n-1}}P(X_{1}>x),\ x>r_{1}.$
By integration by parts,
$\int_{x}^{\infty}\exp(-t^{\beta})dt=\frac{\exp(-x^{\beta})}{\beta}x^{1-\beta}-\frac{\beta-1}{\beta}\int_{x}^{\infty}t^{-\beta}\exp(-t^{\beta})dt$
$\geq\frac{\exp(-x^{\beta})}{\beta}x^{1-\beta}-\frac{\beta-1}{\beta x^{\beta}}\int_{x}^{\infty}\exp(-t^{\beta})dt.$
Hence, there exists $r_{2}>0$ such that for every $x\geq r_{2}$,
$\int_{x}^{\infty}\exp(-t^{\beta})dt\geq\frac{\exp(-x^{\beta})}{2\beta x^{\beta-1}}.$
Now we obtain that
(2.3) $P(M_{n}>x)\geq\frac{nc_{1}}{2^{n}\beta}\frac{\exp(-x^{\beta})}{x^{\beta-1}},\ x>r_{2}.$
We give an upper bound for $P(S_{n}>x)$.
It holds that
$P(S_{n}>x)=c_{1}^{n}\int_{D_{3}}\exp\left(-\sum_{i=1}^{n}x_{i}^{\beta}\right)dx_{1}\cdots dx_{n},$
where $D_{3}\coloneqq\{(x_{1},\dots,x_{n})\in[0,\infty)\colon x_{1}+\cdots+x_{n}>x\}$.
Since $\beta\geq 1$, by convexity, we see that $\sum_{i=1}^{n}x_{i}^{\beta}\geq n^{1-\beta}\left(\sum_{i=1}^{n}x_{i}\right)^{\beta}$. By the change-of-variable $u_{1}=\sum_{i=1}^{n}x_{i},u_{i}=x_{i},2\leq i\leq n$, we obtain that
$\int_{D_{3}}\exp\left(-\sum_{i=1}^{n}x_{i}^{\beta}\right)dx_{1}\cdots dx_{n}\leq\int_{D_{3}}\exp\left(-n^{1-\beta}\left(\sum_{i=1}^{n}x_{i}\right)^{\beta}\right)dx_{1}\cdots dx_{n}$
$\int_{D_{4}}\exp\left(-n^{1-\beta}u_{1}^{\beta}\right)du_{1}\cdots du_{n}=\frac{1}{(n-1)!}\int_{x}^{\infty}u^{n-1}\exp\left(-n^{1-\beta}u^{\beta}\right)du,$
where $D_{4}\coloneqq\{(u_{1},\dots,u_{n})\in[0,\infty)^{n}\colon u_{1}>x,u_{1}>u_{2}+\cdots+u_{n}\}$. By integration by parts,
$\int_{x}^{\infty}u^{n-1}\exp\left(-n^{1-\beta}u^{\beta}\right)du=\int_{x}^{\infty}u^{n-\beta}\left(-\frac{n^{\beta-1}}{\beta}\exp\left(-n^{1-\beta}u^{\beta}\right)\right)^{\prime}du$
$\frac{n^{\beta-1}}{\beta}x^{n-\beta}\exp\left(-n^{1-\beta}x^{\beta}\right)+\frac{(n-\beta)n^{\beta-1}}{\beta}\int_{x}^{\infty}u^{n-\beta-1}\exp\left(-n^{1-\beta}u^{\beta}\right)du.$
Since $u^{n-\beta-1}\leq x^{-\beta}u^{n-1}$ for $u\geq x$, we obtain that
$\int_{x}^{\infty}u^{n-1}\exp\left(-n^{1-\beta}u^{\beta}\right)du$
$\leq\frac{n^{\beta-1}}{\beta}\left(x^{n-\beta}\exp\left(-n^{1-\beta}x^{\beta}\right)+\frac{n-\beta}{x^{\beta}}\int_{x}^{\infty}u^{n-1}\exp\left(-n^{1-\beta}u^{\beta}\right)du\right).$
Hence, there exists $r_{3}>0$ such that for every $x\geq r_{3}$,
$\int_{x}^{\infty}u^{n-1}\exp\left(-n^{1-\beta}u^{\beta}\right)du\leq 2\frac{n^{\beta-1}}{\beta}x^{n-\beta}\exp\left(-n^{1-\beta}x^{\beta}\right).$
Now we obtain that
(2.4) $P(S_{n}>x)\leq 2c_{1}^{n}\frac{n^{\beta-1}}{\beta(n-1)!}x^{n-\beta}\exp\left(-n^{1-\beta}x^{\beta}\right),\ x>r_{3}.$
By (1.3), it holds that $(n!)^{\beta/n}>n^{\beta-1}$. By this, (2.3) and (2.4), we obtain that
$\lim_{x\to\infty}\frac{P(M_{n}>x/(n!)^{1/n})}{P(S_{n}>x)}=\infty.$
This contradicts $S_{n}=(n!)^{1/n}M_{n}$ in distribution. The proof of Theorem 1.1 in the case where $\beta\geq 1$ is completed.
We secondly consider the case where $0<\beta<1$. We recall that $X$ has a subexponential distribution if $\lim_{x\to\infty}\frac{P(X+Y>x)}{P(X>x)}=2$ where $Y$ is an independent copy of $X$. See [3, Section 1.3] for details.
We show that $X_{1}$ has a subexponential distribution. We confirm condition (ii) in [2, Proposition 11]. Since
(2.5) $\log f(x)=x^{\beta}-\log c_{1},\quad x>0,$
there exists $r_{4}>0$ such that for every $x>r_{4}$,
$\frac{d^{2}}{dx^{2}}\left(-\log f(x)\right)=\beta(\beta-1)x^{\beta-2}<0,\quad x>r_{4}.$
In particular, $-\log f(x)$ is concave on $x>r_{4}$. Let $h(x)\coloneqq(\log x)^{2/\beta}$ for $x>1$. Then $\lim_{x\to\infty}xf(h(x))=0$.
By (2.5) and the assumption that $\beta<1$ and the mean value theorem,
$\lim_{x\to\infty}\sup_{|t|\leq h(x)}\left|\log\frac{f(x+t)}{f(x)}\right|=\lim_{x\to\infty}\sup_{|t|\leq h(x)}\left|(x+t)^{\beta}-x^{\beta}\right|=0.$
Therefore, [2, Proposition 11] implies that $X_{1}$ has a subexponential distribution.
By [3, Eq.(1.28)], it holds that if $X_{1}$ has a subexponential distribution, then
(2.6) $\lim_{x\to\infty}\frac{P(M_{n}>x)}{P(S_{n}>x)}=1.$
By Lemma 2.1, if there exists a constant $C$ such that $S_{n}=CM_{n}$ in distribution, then $C=(n!)^{1/n}$. By (2.6) and the fact that
$\lim_{x\to\infty}\frac{P(M_{n}>x)}{P(X_{1}>x)}=n,$
we obtain that
(2.7) $\lim_{x\to\infty}\frac{P(X_{1}>x)}{P(X_{1}>x/(n!)^{1/n})}=\lim_{x\to\infty}\frac{P(M_{n}>x)}{P(M_{n}>x/(n!)^{1/n})}=1.$
By l’Hopital’s rule,
$\lim_{x\to\infty}\frac{P(X_{1}>x)}{x^{1-\beta}\exp(-x^{\beta})}=\frac{c_{1}}{\beta}>0.$
Since $(n!)^{1/n}>1$ for $n\geq 2$,
$\lim_{x\to\infty}\frac{P(X_{1}>x)}{P(X_{1}>x/(n!)^{1/n})}=0.$
This contradicts (2.7). The proof of Theorem 1.1 in the case where $0<\beta<1$ is completed.
Remark $2.$
Set $a_{n}\coloneqq\frac{n\log n}{n\log n-\log(n!)}$ for $n\geq 2$. We can show that $(a_{n})_{n}$ is strictly increasing. Therefore, there exists $N(\beta)\geq 2$ such that (1.3) holds if and only if $n\geq N(\beta)$.
Let $b_{n}\coloneqq\frac{\log(n!)}{n\log n}$. It suffices to show that $(b_{n})_{n}$ is strictly increasing. Let $c_{n}\coloneqq\log(n+1)$ and $d_{n}\coloneqq(n+1)\log(n+1)-n\log n$ for $n\geq 1$. Since $b_{n}=\frac{\sum_{i=1}^{n-1}c_{i}}{\sum_{i=1}^{n-1}d_{i}}$, it suffices to show that $\left(\frac{c_{n}}{d_{n}}\right)_{n}$ is strictly increasing. Let $p(x)\coloneqq\frac{\log(x+1)}{(x+1)\log(x+1)-x\log x}$. Then
$p^{\prime}(x)=\frac{\log(x+1)-((x+1)\log(x+1)-x)\log(1+1/x)}{(x+1)((x+1)\log(x+1)-x\log x)^{2}}.$
By the inequality $\log(1+z)\leq z$ for every $z>0$, it follows that $p^{\prime}(x)>0$ and hence $p(x)$ is increasing.
Remark 2.3.
Let $n=3$. For the half-normal distribution, we have an alternative proof. It is easy to see that
$E\left[S_{3}^{2}\right]=E[(X_{1}+X_{2}+X_{3})^{2}]=3E[X_{1}^{2}]+6E[X_{1}]^{2}=3+\frac{12}{\pi}.$
We also see that
Lemma 2.4.
$E\left[M_{3}^{2}\right]=1+\frac{2\sqrt{3}}{\pi}.$
Proof.
Let $f$ and $F$ be the density and distribution functions of $X_{1}$ respectively. Then the density function of $M_{3}$ is given by $g(x)=3f(x)F(x)^{2}$. Since $xf(x)=-f^{\prime}(x)$, it holds that
$E\left[M_{3}^{2}\right]=3\int_{0}^{\infty}x^{2}f(x)F(x)^{2}dx=-3\int_{0}^{\infty}xF(x)^{2}f^{\prime}(x)dx.$
By integration by parts,
$\int_{0}^{\infty}xF(x)^{2}f^{\prime}(x)dx=-\int_{0}^{\infty}(F(x)^{2}+2xf(x)F(x))f(x)dx.$
Since $2xf(x)^{2}=-\frac{d}{dx}f(x)^{2}$, we obtain that
$\int_{0}^{\infty}(F(x)^{2}+2xf(x)F(x))f(x)dx=\frac{1}{3}-\int_{0}^{\infty}F(x)(f(x)^{2})^{\prime}dx=\frac{1}{3}+\int_{0}^{\infty}f(x)^{3}dx.$
Hence,
$E\left[M_{3}^{2}\right]=1+3\int_{0}^{\infty}f(x)^{3}dx=1+\frac{2\sqrt{3}}{\pi}.$
By Lemma 2.1, if there exists a constant $C$ such that $S_{3}=CM_{3}$ in distribution, then $C$ would be equal to $6^{1/3}$ and hence $3+\frac{12}{\pi}=6^{\frac{2}{3}}\left(1+\frac{2\sqrt{3}}{\pi}\right)$. Then $\pi\in\mathbb{Q}(\sqrt{3},6^{1/3})$, which contradicts the fact that $\pi$ is a transcendental number. We see that $\frac{3+\frac{12}{\pi}}{1+\frac{2\sqrt{3}}{\pi}}$ is approximately $3.24$ and $6^{\frac{2}{3}}$ is approximately $3.30$; they are numerically close to each other.
Acknowledgments The author is supported by JSPS KAKENHI JP22K13928.